# A variational formulation of stochastic thermodynamics. Part I: Finite-dimensional systems

### v1

@@ III.5 Mechanical open system
The dissipative mechanical forces are interpreted as in Sec. III.3. We denote the thermal displacement by $\Gamma$, and the corresponding external entropy flux by $\mathcal{J}_{s,h}(\mathbf{x},\mathbf{p},s)$, with its stochastic counterpart given by $\kappa(\mathbf{x},\mathbf{p},s)\circ\eta(t)$. The external thermodynamic affinity at the boundary is represented by $T^{h}$, the temperature of the external reservoir. This should not be confused with the case of an open thermodynamic system with prescribed fluxes. Additionally, we incorporate the entropy pumping due to external manipulation, denoted by $\mathcal{I}_{p}$, as defined in Eq. (39).
We restrict to 1-dimension for convenience, giving the EoMs:
$p=m\dot{x},$ (83)
@@
$\dot{s}_{f}=\dot{s}-\dot{\Sigma}=\mathcal{J}_{s,h}+\kappa\circ\eta+\mathcal{I}_{p}.$
Since the system is open, the medium EPR $\dot{\Sigma}$ includes an additional contribution to the thermodynamic entropy, accounting for entropy exchanges with sources outside the system $\Omega$. We denote this contribution by $\dot{s}_{f}$, representing the entropy flow entering or leaving the system. See Fig. 3 for an illustration of the system described by Eqs. (LABEL:eq:EoM_open).
@@
Here, $P^{ext}_{W},P^{ext}_{H}$ denote the external power input due to mechanical work and heat transfer, respectively. Equation (III.5) represents the stochastic extension of its deterministic counterpart, as given in Eq. (22) of [51].
Denoting the mechanical and thermal dissipative currents as
@@
the average system EPR gives
$\dot{\mathcal{S}}=$ $\int\frac{d\Omega}{\mathcal{P}}\,\left(\frac{{j}_{d}^{2}}{g^{2}}+\frac{j_{th}^{2}}{\kappa^{2}}\right)+\langle\partial_{p}{\rm f}_{+}\rangle$ (86)
@@
$-\int d\Omega\,\frac{j_{th}}{\kappa^{2}}\left[\mathcal{J}_{s,h}-\kappa\partial_{s}\left(\frac{T^{h}}{T}\kappa\right)\right].$
Given that $\mathcal{I}_{p}=\partial_{p}{\rm f}_{+}$, the average medium EPR yields
@@
Gathering both contributions and enforcing the second law the FDRs are obtained
$\dot{S}_{tot}=\dot{\mathcal{S}}+\langle\!\langle\dot{\Sigma}\rangle\!\rangle=\int\frac{d\Omega}{\mathcal{P}}\,\left(\frac{{j}_{d}^{2}}{D}+\frac{j_{th}^{2}}{\kappa^{2}}\right)\geq 0\iff$ (88)
$\begin{cases}\vskip 8.53581pt\displaystyle-\gamma\frac{p}{m}=-\frac{D}{T}\frac{p}{m}+g\partial_{p}g-\frac{p}{m}g\partial_{s}(g/T),\\ \vskip 8.53581pt\displaystyle\mathcal{J}_{s,h}=\kappa^{2}\left(\frac{T^{h}-T}{T}\right)+\kappa\partial_{s}\left(\frac{T^{h}}{T}\kappa\right).\end{cases}$ (89)
The FDR for the mechanical DoFs coincides with that of a closed mechanical system discussed in Sec. III.3, see Eq. (49). This is because, in the present example, we have not included cross-effects between the dissipative force and the external heat flux, although such couplings could, in principle, be incorporated. Additionally, the FDR associated with the external heat flux resembles that of interconnected systems introduced in Sec. III.4, see Eq. (72). In the macroscopic limit, it again converges to the results of nonequilibrium thermodynamics, i.e., the FDR (89) is consistent with its macroscopic counterpart given by [51]
@@
However, the implications of these FDRs extend further. By applying the path integral (PI) formulation to the system described by Eqs. (LABEL:eq:EoM_open), one can compute the ratio of probabilities for forward and backward trajectories in the thermodynamic phase space $\Omega$. The inverse of the covariance matrix is of the form
@@
with the noise-induced drift $\nu_{p}=g\partial_{p}g-\frac{p}{m}g\partial_{s}(g/T)$. We denote $F^{p},F^{s}$ as the deterministic forces in the momentum and entropy Eqs. (LABEL:eq:EoM_open). Decomposing $D_{\mu\nu}$ into the symmetric and anti-symmetric contributions $D_{\mu\nu}^{+},D_{\mu\nu}^{-}$, respectively, the OM action and its time reversal are given by
@@
where $\mathbf{u}=(x,p,s)$. The ratio of logarithms (11b) then gives
@@
$\left.+\frac{1}{\kappa^{2}}(\mathcal{J}_{s,h}+\kappa\circ\eta)(\mathcal{J}_{s,h}-\kappa\partial_{s}(\kappa T^{h}/T))-\partial_{p}F_{+}^{p}\right),$ (95)
where we denoted the medium entropy defined through the PI in the extended phase space $\Omega$ as $\mathfrak{s}_{m}$, in order to distinguish it from $s_{m}$ in the restricted phase space of Sec. II.2. Differentiating and enforcing the FDRs (89), we find
@@
which comparing with Eqs. (LABEL:eq:EoM_open) gives $\dot{\mathfrak{s}}_{m}=\dot{\Sigma}$. This result is remarkable for a number of reasons. Firstly, it verifies the consistency of our interpretation of $\dot{\Sigma}$ as the medium EPR. Additionally, it shows how the variational approach, specifically the FDRs, provides a well-defined physical meaning to $\dot{\mathfrak{s}}_{m}$, by satisfying the generalized Clausius relation (33). Moreover, this result implies that $\Delta s_{tot}=\Delta s_{sys}+\Delta\Sigma$, with
@@
is compatible, or can be derived, through the KL divergence (9), when the system is constrained by the FDRs. As a consequence, we can prove the following relation, for arbitrary (normalized) initial and final distributions $\mathcal{P}_{0},\mathcal{P}_{t}$ [1]:
@@
where the last equality follows due to normalization, analogously as for the IFT, see (10) in Sec. II.2. The relation (III.5), referred to as the master fluctuation theorem (master FT), serves as a unifying result from which many known FTs in ST can be derived as special cases; see Ref. [1] for a detailed discussion. This demonstrates that a significant portion of the ST literature is encompassed within the variational framework presented here, rather than constituting an independent framework. Notably, the resulting FTs remain valid even in the presence of nonlinear couplings between thermal and mechanical DoFs, thereby broadening the domain of applicability of current formulations. Full derivations supporting this section are provided in Appendix D.
A technical remark is in order: when applying the same computation to systems lacking external irreversible processes, such as in the example of Sec. III.3, one often encounters a singular covariance matrix due to degeneracies arising from energy conservation. In these cases, the master FT can still be recovered by employing the Moore–Penrose pseudo-inverse [60], as illustrated in Appendix E. Consequently, the above results remain valid for closed systems as well.
@@ IV Discussion
Through the examples presented, we have illustrated the application of the variational formulation of stochastic thermodynamics and shown how, by construction, enforcing the second law yields a modeling framework that is fully thermodynamically consistent in the sense defined in Sec. II.4. Specifically, requiring the non-negativity of the average total entropy production rate ensures that two key conditions are satisfied: the thermodynamic entropy is well-defined through a generalized Clausius relation; and for isolated systems, the Fokker–Planck equation admits an equilibrium steady-state distribution consistent with detailed balance. These properties are not imposed ad hoc but instead emerge naturally from the variational construction.
@@
Furthermore, although derived from a geometrically motivated ansatz, the dynamical equations governing entropy production—including those at the level of individual trajectories—retain the same form as in conventional stochastic thermodynamics, up to a reformulation in an extended phase space. Remarkably, this formulation has been shown to coincide with the Kullback–Leibler divergence approach, thereby ensuring consistency with fluctuation theorems and compatibility with the broader theoretical landscape.
In Secs. III.3 and III.4, we showed that the equilibrium distribution arises as a special case of a general equilibrium functional, which recovers classical results—such as the Maxwell-Boltzmann distribution—in appropriate limits. For mechanical systems, the corresponding FDRs generalize the standard formulations by enabling a new class of couplings between thermal and mechanical DoFs, in which the system parameters explicitly depend on the heat bath macrostate $s$. In the free energy variational framework of Ref. [61], it is shown how a Legendre transform of the action shifts this entropy dependence into an explicit temperature dependence of the parameters, allowing systematic inclusion of cases where phenomenological coefficients—such as viscosity—vary with temperature. Such couplings become particularly relevant in systems where this temperature dependence is non-negligible, as demonstrated in e.g. Ref. [62].
@@
A further advantage of this approach is its capacity to model complex systems as interconnections of primitive subsystems. As demonstrated in Sec. III.4, this modular construction naturally yields FDRs for coupled mass and heat transfer that define symmetric, positive-definite matrices, analogous to Onsager’s matrix [57]. In the presence of cross-correlated noise, these matrices consistently capture cross-effects in accordance with nonequilibrium thermodynamic theory.
The variational formulation also extends to open systems operating far from equilibrium. In the macroscopic limit, the obtained FDRs are consistent with phenomenological results from classical nonequilibrium thermodynamics [51, 7]. Importantly, this formalism allows model parameters—such as temperature, transport coefficients or friction—to be state-dependent in the complete phase space $\Omega$, thus offering significant flexibility. Nevertheless, it is important to remark the limitation of this construction which relies on a Langevin description, and therefore assumes a separation of time scales between the slow dynamics of interest and the fast DoFs that have been integrated out.
@@
Furthermore, when multiplicative noise is present, the noise-induced drift terms lead to FDRs that take the form of partial differential equations rather than algebraic constraints. While closed-form solutions may be obtainable in specific cases—as in Ref. [52] for Coulomb friction—this is not guaranteed in general. In such cases, the model can be systematically extended by incorporating these drifts into an effective dissipative force (e.g., ${\rm f}_{-}\rightarrow{\rm f}_{-}+\nu_{p}$), resulting in an algebraic FDR. This strategy has been employed in various contexts [14, 5, 6], and fits naturally within the variational formulation developed here.
Finally, when applying this framework to models that omit relevant energy sources, inconsistencies manifest in the FDR. For example, in systems exhibiting negative friction, the FDR yields $g^{2}/T=\gamma<0$, which is incompatible with $g^{2}\geq 0$ unless one assumes a negative temperature $T<0$. This apparent violation can be resolved by recognizing that such effective energy injection arises from the coupling to an additional degree of freedom.

### v2

@@ III.6 Mechanical open system
The dissipative mechanical forces are interpreted as in Sec. III.4. We denote the thermal displacement by $\Gamma$, and the corresponding external entropy flux by $\mathcal{J}_{s,h}(\mathbf{x},\mathbf{p},s)$, with its stochastic counterpart given by $\kappa(\mathbf{x},\mathbf{p},s)\circ\eta(t)$. The external thermodynamic affinity at the boundary is represented by $T^{h}$, the temperature of the external reservoir. This should not be confused with the case of an open thermodynamic system with prescribed fluxes. Additionally, we incorporate the entropy pumping due to external manipulation, denoted by $\mathcal{I}_{p}$, as defined in Eq. (40).
We restrict to 1-dimension for convenience, giving the EoMs (see Appendix F):
$p=m\dot{x},\quad\dot{\Gamma}=-\frac{\partial L}{\partial s}=T,$ (92)
@@
$\dot{s}_{f}=\dot{s}-\dot{\Sigma}=\mathcal{J}_{s,h}+\kappa\circ\eta+\mathcal{I}_{p},$
where $T$ represents the internal system temperature, to be distinguished from the external bath temperature $T^{h}$. Since the system is open, the medium EPR $\dot{\Sigma}$ includes an additional contribution to the thermodynamic entropy, accounting for entropy exchanges with sources outside the system $\Omega$. We denote this contribution by $\dot{s}_{f}$, representing the entropy flow entering or leaving the system. See Fig. 3 for an illustration of the system described by Eqs. (LABEL:eq:EoM_open).
@@
Here, $P^{ext}_{W},P^{ext}_{H}$ denote the external power input due to mechanical work and heat transfer, respectively. Equation (III.6) represents the stochastic extension of its deterministic counterpart, as given in Eq. (22) of [42].
We denote the ensemble currents associated with the mechanical and thermal irreversible fluxes as
@@
which define the FP probability currents $j_{p},j_{s}$ shown in Appendix F. The physical intuition behind Eqs. (94) is illustrated via the averaged first law:
$\frac{d}{dt}\langle E\rangle$ $=-\int d\Omega\;E\;\nabla_{\Omega}\cdot\bm{j}=\int d\Omega\;\bm{j}\cdot\nabla_{\Omega}E$ (95)
$=\int d\Omega\;\left({\rm f}_{+}\cdot\dfrac{p}{m}\mathcal{P}_{t}+j_{d}\cdot\dfrac{p}{m}+Tj_{s}\right)$
$=\langle{\rm f}_{+}\cdot\dfrac{p}{m}\rangle+\int d\Omega\;T^{h}j_{th}.$
It can be seen how Eq. (95) is consistent with the average of Eq. (III.6), and $j_{th}$ corresponds to the ensemble description of the physical entropy flux $(\mathcal{J}_{s,h}+\kappa\circ\eta)$.
The average system EPR gives (see Appendix F)
$\dot{\mathcal{S}}=$ $\underbrace{\int\frac{d\Omega}{\mathcal{P}}\,\left(\frac{{j}_{d}^{2}}{g^{2}}+\frac{j_{th}^{2}}{\kappa^{2}}\right)}_{\langle\!\langle\dot{\mathfrak{s}}_{\rm tot}\rangle\!\rangle}+\langle\partial_{p}{\rm f}_{+}\rangle$ (96)
@@
$-\int d\Omega\,\frac{j_{th}}{\kappa^{2}}\left[\mathcal{J}_{s,h}-\kappa\partial_{s}\left(\frac{T^{h}}{T}\kappa\right)\right]$
$=$ $\langle\!\langle\dot{\mathfrak{s}}_{\rm tot}\rangle\!\rangle-\langle\!\langle\dot{\mathfrak{s}}_{m}\rangle\!\rangle.$
The identification of the corresponding terms with $\langle\!\langle\dot{\mathfrak{s}}_{\rm tot}\rangle\!\rangle$ and $\langle\!\langle\dot{\mathfrak{s}}_{m}\rangle\!\rangle$ follows from a direct verification, see Eq. (III.6) below. Given that $\mathcal{I}_{p}=\partial_{p}{\rm f}_{+}$, the average medium EPR yields
@@
Adding both contributions the total average EPR gives
$\dot{S}_{\rm tot}$ $=\dot{\mathcal{S}}+\langle\!\langle\dot{\Sigma}\rangle\!\rangle$ (98)
Following the general argument in Appendix B, enforcing the second law requires that the terms in braces should vanish, giving the FDRs and associated total EPR:
$\begin{cases}\vskip 8.53581pt\displaystyle-\gamma\frac{p}{m}=-\frac{D}{T}\frac{p}{m}+g\partial_{p}g-\frac{p}{m}g\partial_{s}(g/T),\\ \vskip 8.53581pt\displaystyle\mathcal{J}_{s,h}=\kappa^{2}\left(\frac{T^{h}-T}{T}\right)+\kappa\partial_{s}\left(\frac{T^{h}}{T}\kappa\right),\end{cases}$ (99)
$\Rightarrow\dot{S}_{\rm tot}=\int\frac{d\Omega}{\mathcal{P}}\,\left(\frac{{j}_{d}^{2}}{D}+\frac{j_{th}^{2}}{\kappa^{2}}\right)\geq 0.$ (100)
The FDR for the mechanical DoFs coincides with that of a closed mechanical system discussed in Sec. III.4, see Eq. (54). This is because, in the present example, we have not included cross-effects between the dissipative force and the external heat flux, although such couplings could, in principle, be incorporated. Additionally, the FDR associated with the external heat flux resembles that of interconnected systems introduced in Sec. III.5, see Eq. (81), corresponding to the discrete Fourier’s law. In the macroscopic limit, it again converges to the results of nonequilibrium thermodynamics, i.e., the FDR (99) is consistent with its macroscopic counterpart given by [42]
@@
In Sec. III.3 we anticipated that the FDRs obtained through the second law axiom restore LDB by ensuring the identity $\dot{\mathfrak{s}}_{m}=\dot{\Sigma}$, and we further demonstrated this result throughout Secs. III.4-III.5. We now confirm this result in the far from equilibrium setting by providing an explicit demonstration and examining its implications.
Applying the PI formulation to the system described by Eqs. (LABEL:eq:EoM_open), we compute the ratio of probabilities for forward and backward trajectories in the thermodynamic phase space $\Omega$. The inverse of the covariance matrix is of the form
@@
with the noise-induced drift $\nu_{p}=g\partial_{p}g-\frac{p}{m}g\partial_{s}(g/T)$. We denote $F^{p}=-\partial_{x}U-\gamma p/m+{\rm f}_{+},\ F^{s}=\frac{\gamma}{T}(p/m)^{2}+\frac{T^{h}}{T}\mathcal{J}_{s,h}$ as the deterministic forces in the momentum and entropy Eqs. (LABEL:eq:EoM_open). Decomposing $D_{\mu\nu}$ into the symmetric and anti-symmetric contributions $D_{\mu\nu}^{+},D_{\mu\nu}^{-}$, respectively, the OM action and its time reversal are given by
@@
where $\mathbf{u}=(x,p,s)$. The log ratio (11b) then gives
@@
$\left.+\frac{1}{\kappa^{2}}(\mathcal{J}_{s,h}+\kappa\circ\eta)(\mathcal{J}_{s,h}-\kappa\partial_{s}(\kappa T^{h}/T))-\partial_{p}F_{+}^{p}\right).$ (106)
Differentiating and enforcing the FDRs (99), we find
@@
which comparing with Eqs. (LABEL:eq:EoM_open) gives the identity $\dot{\mathfrak{s}}_{m}=\dot{\Sigma}$. Crucially, the variational formulation distinguishes the medium EPR $\dot{\Sigma}$ from the thermodynamic EPR $\dot{s}$ by properly accounting for external entropy flows. As a result, it correctly identifies $\dot{\mathfrak{s}}_{m}=\dot{\Sigma}$, rather than $\dot{\mathfrak{s}}_{m}=\dot{s}$.
This verification establishes the key results anticipated in Sec. III.3 in the far from equilibrium regime: (i) it confirms the interpretation of $\dot{\Sigma}$ as the medium EPR, and of the FDRs as the link between the second law axiom and LDB [Eq. (45)]; (ii) it shows that the FDRs provide a well-defined physical meaning to $\dot{\mathfrak{s}}_{m}$ by ensuring the generalized Clausius relation (33); (iii) it demonstrates that, under the FDRs, the decomposition $\Delta s_{\rm tot}=\Delta s_{sys}+\Delta\Sigma$, with
@@
is consistent with the KL divergence representation of entropy production (41), i.e. $\langle\Delta s_{\rm tot}\rangle=\langle\Delta\mathfrak{s}_{\rm tot}\rangle$.
As a consequence, we can prove the following relation, for arbitrary (normalized) initial and final distributions $\mathcal{P}_{0},\mathcal{P}_{t}$ [79]:
@@
where the last equality follows due to normalization, analogously as for the IFT, see (10) in Sec. II.2. The relation (III.6), referred to as the master fluctuation theorem (master FT), serves as a unifying result from which many known FTs in ST can be derived as special cases; see Ref. [79] for a detailed discussion. This demonstrates that a significant portion of the ST literature is encompassed within the variational framework presented here, rather than constituting an independent framework.
Notably, the resulting FTs remain valid even in the presence of nonlinear couplings between thermal and mechanical DoFs, thereby broadening the domain of applicability of current formulations. Moreover, the meaning of the master FT is elevated: in our framework the quantity appearing in the exponent, $\Delta\Sigma$, acquires a clear physical interpretation as the medium entropy change, rather than being understood solely as an information-theoretic term. Ultimately, Eq. (III.6) is a direct consequence of LDB [Eq. (45)] in the extended phase space, and thus applies to all examples considered in this work.
A technical remark is in order: when applying the same computation to systems lacking external irreversible processes, such as in the example of Sec. III.4, one often encounters a singular covariance matrix due to degeneracies arising from energy conservation. In these cases, the master FT can still be recovered by employing the Moore–Penrose pseudo-inverse [67], as illustrated in Appendix D.
@@ IV Discussion
Through the examples presented above, we have demonstrated how the variational formulation of ST provides a systematic route to thermodynamically consistent modeling. In particular, enforcing the second law axiom within the variational principle ensures consistency in the precise sense defined in Sec. II.4, namely, compatibility with LDB. Imposing the non-negativity of the average total EPR directly yields the corresponding FDRs, or equivalently, the LDB condition Eq. (45), and thus, the thermodynamic entropy is well defined through the Clausius relation and, for isolated systems, the associated FP equation admits an equilibrium steady state. Crucially, none of these properties is imposed ad hoc: they arise naturally from the structure of the variational construction itself.
@@
Although Appendix B provides a formal argument for the emergence of FDRs as solutions of the second law inequality within the variational framework, and their consistency is verified throughout Secs. III.4III.6, a rigorous proof of uniqueness remains an open problem. This observation points to a promising direction for future work: elucidating the general correspondence between the second law axiom and LDB, and clarifying its relation to alternative constraints, such as the orthogonality condition arising in generalized Langevin equations [49]. Moreover, although derived from a geometrically motivated ansatz, the dynamical equations governing entropy production—including those at the level of individual trajectories—retain the same form as in conventional ST, up to a reformulation in an extended phase space. Remarkably, once the FDRs are specified, this formulation coincides with the KL–divergence approach, thereby ensuring consistency with fluctuation theorems and compatibility with the broader theoretical framework.
In Secs. III.4 and III.5, it is shown that the equilibrium distribution arises as a special case of a general equilibrium functional (59), which recovers classical results—namely, the MB distribution—in appropriate limits. A natural direction for future work is to explore equilibrium distributions in more general settings beyond the MB case and to compare the resulting predictions with experimental observations, for instance in coupled system–bath dynamics.
The derived FDRs generalize the standard formulations by allowing a new class of couplings between thermal and configurational DoFs. In the examples discussed in Secs. III.4III.6, these couplings represent system–bath interactions, whereas in Sec. III.5 they correspond to bath–bath interactions. Within this framework, thermodynamic affinities such as temperature and chemical potential remain well defined, and the bath dynamics themselves follow from the same variational principle. Hence, the underlying mathematical structure remains consistent even in regimes where LDB and the definition of thermodynamic affinities are typically difficult to guarantee [9, 61].
This is due to the fact that the thermodynamic phase space accounts for the complete super-system in a top-down fashion, analogous to the microcanonical formulation [93], in contrast to conventional ST formulations where the bath is not explicitly included, leading to ambiguities and requiring heat and entropy to be inferred indirectly from the system dynamics. In the bottom-up approach of Ref. [93], developed under the same fundamental assumptions as here, it is shown that explicitly accounting for the full super-system allows ST to be extended to strongly-coupled systems and tunable system–bath interactions. Crucially, the LDB condition remains unchanged, and the resulting theory retains the same structural form as in the weak-coupling regime [92], suggesting that the present approach should be tested in such regimes as well. Another shared feature is that LDB can be derived without recourse to path probabilities.
Nonetheless, the physical meaning and range of applicability of this extended description remain open questions, indicating an interesting direction for future research involving both experimental and theoretical validation. For instance, in the free energy variational framework of Ref. [43], it is shown that a Legendre transform of the action shifts this entropy dependence into an explicit temperature dependence of the parameters, thereby allowing systematic inclusion of cases where phenomenological coefficients—such as viscosity—vary with temperature. Such couplings become particularly relevant in systems where this temperature dependence is non-negligible, as demonstrated, for example, in Ref. [26].
@@
A further advantage of this approach is its capacity to model complex systems as interconnections of primitive subsystems. As demonstrated in Sec. III.5, this modular construction naturally yields FDRs for coupled mass and heat transfer that define symmetric, PD matrices consistent with Onsager symmetry [70]. In the presence of cross-correlated noise, these matrices consistently capture cross-effects in accordance with nonequilibrium thermodynamic theory and LDB.
The variational formulation also extends to open systems operating far from equilibrium. In the macroscopic limit, the obtained FDRs are consistent with phenomenological results from classical nonequilibrium thermodynamics [42, 44]. Importantly, this formalism allows model parameters—such as temperature, transport coefficients or friction—to be state-dependent in the complete phase space $\Omega$, thus offering significant flexibility.
@@
Furthermore, in the presence of multiplicative noise, the noise-induced drift terms lead to FDRs that take the form of partial differential equations rather than algebraic constraints. While closed-form solutions may be obtainable in specific cases—as in Ref. [28] for Coulomb friction—this is not guaranteed in general. In such cases, the model can be systematically extended by incorporating these drifts into an effective dissipative force (e.g., ${\rm f}_{-}\rightarrow{\rm f}_{-}+\nu_{p}$), resulting in an algebraic FDR. This strategy has been employed in various contexts [57, 12, 64], and fits naturally within the variational formulation developed here.
Finally, although this framework requires an initial specification of the thermal variables, any inconsistency between this choice and the underlying dynamics will arise as a mathematical inconsistency. For instance, in systems exhibiting negative friction, the FDR implies $g^{2}/T=\gamma<0$, which is incompatible with $g^{2}\geq 0$ unless one assumes a negative temperature $T<0$. However, since temperature is defined independently of $\gamma$, such a condition cannot be imposed. Consequently, LDB breaks down, appearing as an inconsistency in the FDRs. This apparent violation can be resolved by recognizing that an effective energy input arises from coupling to an additional DoF—meaning that the initial model neglected a thermodynamically relevant variable [5]. Hence, the framework does not arbitrarily enforce LDB but instead reveals when a more complete thermodynamic phase space must be considered, thereby guiding the modeling process.